Solve: log(subscript 6)x + log(subscript 6)(x-5) = 2
Answer by edjones(7569)

log(6)x+log(6)(x-5)=2
log(6)(x(x-5))=2
log(6)(x^2-5x)=2
x^2-5x=6^2
x^2-5x-36=0
(x-9)(x+4)=0
x=9 is the only answer because neg. numbers are not in the domain of logs.
Ed